1. A car of mass m = 1,500 kg (~midsized sedan), originally moving at 10 m/s (~ 22 mi/hr), accelerates until it reaches a speed of v = 20 m/s (~ 44 mi/hr). How much kinetic energy has the car gained?

Please explain in detail.

Thank you.
Answer #1

Kinetic energy gained = 1/mv2 - 1/2 mu2 =1/2×1500(400_100) =225000 j